# url]sites.google.com/view/mslee2

### v1

@@ 5.1 Bypassing term
The instability caused by the exponential growth of the $\sinh$ term in the BV equation, particularly when applied to PINNs, can be attributed to the sensitivity of the overpotential $\eta$ to changes in the network parameters. Instead of linearizing the BV equation, we propose a novel approach to reduce the sensitivity of $\eta$ during the optimization process, making it possible to handle the fully non-linear BV equation.
@@
$\widetilde{j}_{i}=\dfrac{2kL^{2}aF}{\sigma^{\mathrm{eff}}\psi_{\mathrm{ref}}}(c_{s,\mathrm{max}}-\widetilde{c}_{s,\mathrm{surf}})^{0.5}\widetilde{c}_{s,\mathrm{surf}}^{0.5}\widetilde{c}^{0.5}\sinh\beta_{\theta},$ (9)
instead of the formulation in (4). Then, an additional penalty term is incorporated into the PINN loss to ensure $\beta_{\theta}$ accurately approximates $(F/2RT)\eta$:
@@
We refer to this strategy as bypassing. The purpose of this bypassing strategy is to mitigate the risk of exploding loss values during training, which arises due to the instability of the BV equation.
In [36], Wang et al. observed that an imbalance between the gradients of the PDE loss ($\nabla_{\theta}\mathcal{L}_{\mathrm{PDE}}$ and the boundary condition loss $\nabla_{\theta}\mathcal{L}_{\mathrm{BC}}$) can lead to a wide range of eigenvalues in the Hessian matrix, resulting in ill-conditioned training dynamics. This emphasizes the importance of maintaining a balanced gradient distribution for effective training. From this point of view, we can verify the effectiveness of the bypassing strategy. By conducting an eigenvalue analysis, we observed a significant reduction in the condition number after introducing the bypassing term. Although memory constraints required us to perform this analysis with a smaller neural network compared to the one used for the full P2D model in Section 6.1, the results still sufficiently demonstrate the effectiveness of the proposed strategy. The results are visualized in Figure 5.
@@ 5.2 Secondary conservation laws
To address this, we introduce additional constraints based on secondary conservation laws to more accurately approximate the reaction rates $\widetilde{j}_{i}$, which, in turn, guide the approximate solutions toward the true solution. Specifically, we enforce the following secondary conservation laws:
@@
For the paradigm example presented in Section 4, we can see that this strategy helps prevent the network from converging to a shifted or incorrect solution. For the paradigm example, the corresponding secondary conservation can be expressed as:
@@
As previously discussed, using a typical PINN loss alone is not sufficient for solving the paradigm example effectively, especially when dealing with the small factor $j_{0}$. However, by incorporating the additional constraints (11) into the PINN loss function, the PINNs can predict the solution more accurately, as shown in Figure 6.
@@
As demonstrated in the paradigm example, by combining secondary conservation laws with the PINN framework, we can mitigate the challenges posed by the BV equation and improve the network’s ability to capture the true solution with greater accuracy, despite the computational challenges.
Figure 7: A schematic diagram of our proposed PINN architecture for the P2D model of a lithium-ion battery cell.
@@ 5.3 Ablation Study
We performed an ablation study to evaluate the individual contributions and combined effectiveness of the methods proposed in this section. The results are summarized in Table 2.
@@
Table 2: Ablation study results showing relative $L^{2}$ errors. B: Results obtained using the bypassing term to approximate the stiffest component of the BV equation. SC: Results obtained using the secondary conservation law to enhance solution accuracy by preventing convergence to trivial solutions. B+SC: Results obtained when both strategies are applied simultaneously.
The Baseline configuration represents the simplest implementation of the PINN framework, without the bypassing term or the secondary conservation law. The B configuration includes the bypassing term to address the stiffness of the BV equation, while the SC configuration incorporates the secondary conservation law into the PINN loss to enhance solution accuracy. The results highlight the challenges posed by the ill-conditioned BV equation in the Baseline and SC configurations, where most line search proposals are rejected, leading to significantly prolonged training times exceeding 8 hours. Notably, while the introduction of the bypassing term (B) improves stability and reduces training time, it is insufficient on its own to prevent failure modes inherent to PINNs. This underscores the importance of the secondary conservation law (SC) as a complementary component to ensure convergence to physically meaningful solutions. Overall, our ablation study demonstrates that neither method is redundant; both the bypassing term and the secondary conservation law are crucial for the efficient and successful simulation of the P2D model. This highlights the robustness of our proposed framework in overcoming the challenges associated with nonlinear BV equations in PINNs.
@@ 6 Numerical Experiments
We performed numerical experiments for both forward and inverse problems of the P2D model to evaluate the effectiveness of our proposed methods. In our numerical experiments, we used a neural network architecture with a width of 32 and a depth of 4. The optimization process followed a two-step approach, combining ADAM [37] and L-BFGS [38] optimizers, implemented in the Optax package [39] and the JAXopt package [40], respectively. Collocation points were dynamically sampled for ADAM but kept fixed during L-BFGS, using $2^{14}$ points at each step. The ADAM optimizer employed a cosine decay learning rate schedule, starting from $10^{-4}$, and the L-BFGS history size was set to 50. Our network utilized the sine activation function, initialized with the SIREN scheme [41]. Computations were carried out on an Nvidia A100 GPU, and the framework JAX [42] was used for automatic differentiation, while visualizations were created using Makie.jl [43].
@@ 6.1 Forward Problem
We present the PINN simulation results for the forward problem of the P2D model, demonstrating the effectiveness of our proposed methods in addressing the challenges posed by the fully nonlinear BV equation (1). The simulations were performed using 100k iterations of the Adam optimizer, followed by 100k iterations of the L-BFGS optimizer. Figures 8 and 9 illustrate the spatial distribution and temporal evolution of key battery state variables, evaluated at $t=3500$ and $x=0$, respectively. Unlike vanilla PINNs, which fail to produce meaningful predictions and deviate significantly from the reference solution, our proposed approach delivers highly accurate results. This close agreement between our PINN solution and the reference solution not only validates the robustness of our computational strategies but also highlights their predictive accuracy over long time horizons.
@@ 6.2 Inverse Problem
The minimizer $\left(u_{\theta^{\star}},\alpha^{\star}\right)$ of this loss function will yield a solution $(u_{\theta^{\star}},\alpha^{\star})$ that satisfies the governing PDE and boundary conditions, while also fitting the observed data closely, thus providing a solution to the inverse problem. With a more reliable method for solving the forward problem now in place, we are prepared to address inverse problems using the reference solution as our data. Our focus is on inverse problems related to battery length, which is considered a crucial factor in the performance of Li-ion batteries [44]. In the following, we perform experiments on three inverse tasks: estimating the battery length, the length ratio, and both battery length and length ratio simultaneously. These tasks are evaluated under two scenarios: the first uses noise-free observational data, while the second introduces 5% Gaussian noise into the observational data. The detailed results are presented in Table 3. Our proposed methodology demonstrates strong predictive performance across all scenarios. Notably, even in the presence of noisy observational data, the predictions remain comparable to those obtained with clean data. This robustness to noise underscores the reliability and effectiveness of our approach in solving inverse problems, even under challenging conditions.
Length Length, Noisy Data Ratio Ratio, Noisy Data
True $L=3.26$ $(L_{p},L_{n})=(1.74,1.00)$
Prediction 3.220 3.221 (1.638, 1.002) (1.638, 1.002)
Error 1.227e-2 1.196e-2 (4.739e-2, 1.827e-3) (4.767e-2, 2.0e-3)
Data Loss 3.913e-3 4.782e-2 2.598e-3 4.648e-2
Table 3: Results for inverse problems. Error is a relative error, $|\mathrm{Prediction}-\mathrm{True}|/|\mathrm{True}|$.
@@
Battery length. The first example involves estimating the total battery length, $L$, while maintaining a fixed ratio of $L_{p}:L_{s}:L_{n}=1.74:0.52:1.0$. Although this problem involves only a single parameter, it is nontrivial due to its impact on 11 coefficients in the P2D model after non-dimensionalization, particularly those associated with spatial partial derivatives along $x$. To parameterize $L$, we expressed it as $\hat{L}\cdot 10^{-4}$ with an initial value of $\hat{L}=1$. After 150,000 iterations of the Adam optimizer, we obtained $\hat{L}=3.22$, corresponding to a relative error of 1.23% compared to the true value $L=3.26$. The entire computation was completed in approximately 30 minutes.
@@
Length ratio. The second example focuses on estimating the ratio of the battery section lengths, $L_{p}:L_{s}:L_{n}$, while keeping the total battery length fixed at $L=3.26\cdot 10^{-4}$. This task introduces additional complexity due to the dynamically changing domain lengths during optimization, which limits the effectiveness of the L-BFGS optimizer. To simplify the parameterization, we used two parameters, $\rho_{p}$ and $\rho_{n}$, since the constraint $\rho_{p}+\rho_{s}+\rho_{n}=1$ reduces the problem to two degrees of freedom. The initial values were set to $\hat{\rho}_{p}=\hat{\rho}_{n}=0.4$. Our method estimated $\hat{L}_{p}=\hat{\rho}_{p}\cdot L=1.64\cdot 10^{-4}$ and $\hat{L}_{n}=\hat{\rho}_{n}\cdot L=1.002\cdot 10^{-4}$, achieving relative errors of 5.75% and 0.2% for the positive and negative electrode lengths, respectively. The computation was completed in approximately 40 minutes.
@@
Battery length and length ratio. Finally, we extended the problem to simultaneously estimate both the battery length and the length ratio. This combined task represents a significantly more challenging problem due to the increased dimensionality and the inherent coupling between the two parameters. Accurate estimation in such scenarios requires the framework to handle dynamic changes in both domain lengths and parameter dependencies throughout the optimization process. Starting with the same initialization for $\hat{L},\hat{\rho}_{p}$, and $\hat{\rho}_{n}$, our method estimated $\hat{L}=2.996$, $\hat{L}_{p}=1.511\cdot 10^{-4}$, and $\hat{L}_{n}=9.401\cdot 10^{-5}$. The relative errors for these estimates were $8.098\cdot 10^{-2}$, $1.316\cdot 10^{-1}$, and $5.99\cdot 10^{-2}$, respectively. These results demonstrate the robustness of our method in tackling complex coupled inverse problems.
@@ 7 Discussion and conclusion
This study has addressed the challenges of simulating lithium-ion batteries using the P2D model in combination with PINNs and has proposed effective strategies to overcome these obstacles. Our methods enable reliable and accurate simulations of the P2D model, incorporating the fully nonlinear Butler-Volmer equation. By developing an enhanced forward solver, we successfully tackled inverse problems related to battery section lengths, achieving remarkable accuracy and computational efficiency. Despite the inherent complexity of these problems, our framework delivers results that hold significant engineering relevance for the design and optimization of lithium-ion batteries. These achievements highlight the practical applicability of our approach and its ability to solve inverse problems that might not be infeasible for traditional numerical methods. However, there is still room for further development. For instance, extending our framework to incorporate multiphysics phenomena such as thermal effects or aging, which are currently excluded from our equations, presents an exciting and challenging avenue for future research. Additionally, building on our forward problem solver to develop an operator network capable of real-time predictions is another promising direction.
@@ Acknowledgements
The work of M.-S. Lee was supported by Basic Science Research Programs through the National Research Foundation of Korea(NRF) funded by the Ministry of Education (RS-2023-00244475 and RS-2024-00462755). The work of Y. Hong was supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF), funded by the Ministry of Education (NRF-2021R1A2C1093579), the Korean government (MSIT) (RS-2023-00219980), and Hyundai Motor Company. The authors also thank the Performance Concept Technology Research Lab at Hyundai Motor Company for their valuable discussions.

### v2

@@ 5.1 Bypassing term
The instability caused by the exponential growth of the $\sinh$ term in the BV equation, particularly when applied to PINNs, can be attributed to the sensitivity of the overpotential $\eta$ to changes in the network parameters. In fact, the PINN approach presented in Section 3 approximates the overpotential $\widetilde{\eta}$ using multiple approximate solutions $\widetilde{\psi}$, $\widetilde{\phi}$, $\widetilde{c}_{s}$, according to the definition of overpotential $\psi-\phi-U(c_{s,\mathrm{surf}}/c_{s,\mathrm{max}})$. In this way, $\widetilde{\eta}$ depends on several networks, making it not straightforward to initialize $\widetilde{\eta}$ at a small value. Consequently, the approximate ionic flux $\widetilde{j}_{i}$ become highly sensitive to errors in any of the associated approximations, which results in an instable training. In addition, because each network is also involved in different types of losses terms, $\widetilde{\eta}$ can be affected by various parts of the overall loss function, leading to difficulties in maintaining a proper approximation. For this, we propose a novel approach to reduce the sensitivity of $\eta$ during the optimization process, making it possible to handle the fully non-linear BV equation.
@@
$\widetilde{j}_{i}=2kc_{s,i,\mathrm{max}}c_{\mathrm{ref}}^{0.5}\left(1-\widetilde{c}_{s,i,\mathrm{surf}}\right)^{0.5}\widetilde{c}_{s,i,\mathrm{surf}}^{0.5}\widetilde{c}_{i}^{0.5}\sinh\left(\beta_{\theta}\right)$ (10)
instead of the formulation in (3). Then, an additional penalty term $\mathcal{L}_{\beta}$ is incorporated into the PINN loss to ensure $\beta_{\theta}$ accurately approximates $(F/2RT)\eta$:
@@
We refer to this strategy as bypassing. By assigning a dedicated single network to approximate the overpotential $\eta$, we isolate the approximation inside the hyperbolic sine function from the other networks and prevent large interference among several types of loss terms. Furthermore, this design facilitates initializing $\beta_{\theta}$ at an $O(1)$ scale, which helps stabilize the early stages of training. Altogether, these mitigate the risk of exploding loss values.
In [37], Wang et al. observed that an imbalance between the gradients of the PDE loss ($\nabla_{\theta}\mathcal{L}_{\mathrm{PDE}}$) and the boundary condition loss ($\nabla_{\theta}\mathcal{L}_{\mathrm{BC}}$) can lead to a wide range of eigenvalues in the Hessian matrix, resulting in ill-conditioned training dynamics. This emphasizes the importance of maintaining a balanced gradient distribution for effective training. From this point of view, we can verify the effectiveness of the bypassing strategy. By conducting an eigenvalue analysis, we observed a significant reduction in the condition number after introducing the bypassing term. Although memory constraints required us to perform this analysis with a smaller neural network compared to the one used for the full P2D model in Section 6.1, the results still sufficiently demonstrate the effectiveness of the proposed strategy. The results are visualized in Figure 6.
@@ 5.2 Secondary conservation laws
To mitigate this discrepancy between loss and accuracy, we introduce additional constraints based on secondary conservation laws to more accurately approximate the reaction rates $\widetilde{j}_{i}$, which, in turn, guide the approximate solutions toward the true solution. Specifically, we enforce the following secondary conservation laws:
@@
For the toy example presented in Section 4, we can see that this strategy helps prevent the network from converging to a shifted or incorrect solution. For the toy example, the corresponding secondary conservation can be expressed as:
@@
As previously discussed, using a typical PINN loss alone is not sufficient for solving the toy example effectively, especially when dealing with the small factor $j_{0}$. However, by incorporating the additional constraints (13) into the PINN loss function, the PINNs can predict the solution more accurately, as shown in Figure 7.
@@
As demonstrated in the toy example, by combining secondary conservation laws with the PINN framework, we can mitigate the challenges posed by the BV equation and improve the network’s ability to capture the true solution with greater accuracy, despite the computational challenges.
Figure 8: A schematic diagram of our proposed PINN architecture for the P2D model of a Li-ion battery cell.
@@ 5.3 Ablation study
We performed an ablation study to evaluate the individual and combined effectiveness of the methods proposed in this section. The results are summarized in Table 2.
@@
Table 2: Ablation study results showing relative $L^{2}$ errors. B: Using the bypassing term to approximate the stiffest component of the BV equation. SC: Using the secondary conservation law to enhance solution accuracy by preventing convergence to trivial solutions. B + SC: Using both strategies simultaneously.
The Baseline configuration represents the simplest implementation of the PINN framework presented in Section 3, without both the bypassing term and the secondary conservation law. The B configuration includes the bypassing term to address the stiffness of the BV equation, whereas the SC configuration incorporates the secondary conservation law into the PINN loss to mitigate the discrepancy between loss and accuracy.
Necessity of bypassing terms
The Baseline and SC configurations both exhibit extremely large losses (on the order of $10^{48}$) and require over 8 hours of training, as neither addresses the high sensitivity of the BV equation. This underscores the need to mitigate the ill-conditioning posed by the nonlinear BV term. By employing the bypassing method, we successfully reduce the PINN loss to $7.32\times 10^{-8}$, indicating that the instability by the BV equation has been effectively resolved. The bypassing method remains effective when combined with the secondary conservation law, achieving a still low loss value of $9.15\times 10^{-6}$.
Necessity of secondary conservation laws
Although the bypassing term ensures faster training and a small loss, the accuracy achieved with the bypassing method alone is still insufficient. In other words, achieving a small value for the PINN loss (4) with bypassing loss (11) does not guarantee accuracy. As noted in Section 5.2, the secondary conservation laws are required to resolve this discrepancy. Indeed, once a sufficiently small loss is achieved through the bypassing method, the secondary conservation plays an important role in resolving the discrepancy between loss and accuracy, improving accuracy to the order of $10^{-2}$ or lower.
Overall, our ablation study shows that neither method is redundant; both the bypassing term and the secondary conservation law are essential for an efficient and accurate simulation of the P2D model. This result underscores the robustness of our proposed framework in overcoming the challenges posed by nonlinear BV equations in PINNs.
@@ 6 Numerical experiments
We performed numerical experiments for both forward and inverse problems of the P2D model to evaluate the effectiveness of our proposed methods. In our numerical experiments, we used a neural network architecture with a width of 32 and a depth of 4. The optimization process followed a two-step approach, combining ADAM [38] and L-BFGS [39] optimizers, implemented in the Optax package [40] and the JAXopt package [41], respectively. Collocation points were dynamically sampled for ADAM but kept fixed during L-BFGS, using $2^{14}$ points at each step. The ADAM optimizer employed a cosine decay learning rate schedule, starting from $10^{-4}$, and the L-BFGS history size was set to 50. Our network utilized the sine activation function, initialized with the SIREN scheme [42]. For reference solution, we used $\Delta t=1$ for time-stepping up to the final time $\tau=3500$, and $N_{x}=50$ for spatial discretization. Computations were carried out on an Nvidia A100 GPU, and the framework JAX [43] was used for automatic differentiation, while visualizations were created using Makie.jl [44].
@@ 6.1 Forward problem
We present the PINN simulation results for the forward problem of the P2D model, demonstrating the effectiveness of our proposed methods in addressing the challenges posed by the fully nonlinear BV equation (1). The simulations were performed using 100k iterations of the Adam optimizer, followed by 100k iterations of the L-BFGS optimizer. Figures 9 and 10 illustrate the spatial distribution and temporal evolution of key battery state variables, evaluated at $t=3500$ and $x=0$, respectively. Unlike vanilla PINNs, which fail to produce meaningful predictions and deviate significantly from the reference solution, our proposed approach delivers accurate results. A includes results for additional experiments with different physical parameters. This close agreement between our PINN solution and the reference solution not only validates the robustness of our computational strategies but also highlights their predictive accuracy over long time horizons.
@@ 6.2 Inverse problem
The minimizer $\left(u_{\theta^{\star}},\alpha^{\star}\right)$ of this loss function will yield a solution $(u_{\theta^{\star}},\alpha^{\star})$ that satisfies the governing PDE and boundary conditions, while also fitting the observed data closely, thus providing a solution to the inverse problem. With a more reliable method for solving the forward problem now in place, we are prepared to address inverse problems using the reference solution as our data. Our focus is on inverse problems related to battery length, which is considered a crucial factor in the performance of Li-ion batteries [45]. In the following, we perform experiments on three inverse tasks: estimating the battery length (Table 3), the length ratio (Table 4), and both battery length and length ratio simultaneously (Table 5). These tasks are evaluated under two scenarios: the first uses noise-free observational data, while the second introduces 5% Gaussian noise into the observational data. Our proposed methodology demonstrates strong predictive performance across all scenarios. Notably, even in the presence of noisy observational data, the predictions remain comparable to those obtained with clean data. This robustness to noise underscores the reliability and effectiveness of our approach in solving inverse problems, even under challenging conditions.
@@
Battery length The first example involves estimating the total battery length, $L$, while maintaining a fixed ratio of $L_{p}:L_{s}:L_{n}=1.74:0.52:1.0$. Although this problem involves only a single parameter, it is nontrivial due to its impact on 11 coefficients in the P2D model after non-dimensionalization, particularly those associated with spatial partial derivatives along $x$. To parameterize $L$, we expressed it as $\hat{L}\cdot 10^{-4}$ with an initial value of $\hat{L}=1$. After 150,000 iterations of the Adam optimizer, we obtained $\hat{L}=3.21$, corresponding to a relative error of 1.51% compared to the true value $L=3.26$. The entire computation was completed in approximately 30 minutes.
Method w/o Noise w/ Noise
Baseline 69.53 69.53
B 44.71 44.85
SC 69.59 69.59
B + SC 1.51 1.58
Table 3: Relative errors (%) for the total length ($L$) estimation problem.
@@
Length ratio The second example focuses on estimating the ratio of the battery section lengths, $L_{p}:L_{s}:L_{n}$, while keeping the total battery length fixed at $L=3.26\cdot 10^{-4}$. This task introduces additional complexity due to the dynamically changing domain lengths during optimization, which limits the effectiveness of the L-BFGS optimizer. To simplify the parameterization, we used two parameters, $\rho_{p}$ and $\rho_{n}$, since the constraint $\rho_{p}+\rho_{s}+\rho_{n}=1$ reduces the problem to two degrees of freedom. The initial values were set to $\hat{\rho}_{p}=\hat{\rho}_{n}=0.4$. The B + SC method estimated $\hat{L}_{p}=\hat{\rho}_{p}\cdot L=1.68\cdot 10^{-4}$ and $\hat{L}_{n}=\hat{\rho}_{n}\cdot L=0.9536\cdot 10^{-4}$, achieving relative errors of 3.21% and 4.64% for the positive and negative electrode lengths, respectively. The computation was completed in approximately 40 minutes.
Method w/o Noise w/ Noise
$L_{p}$ $L_{n}$ $L_{p}$ $L_{n}$
Baseline 26.14 4.65 26.14 4.68
B 2.58 4.33 2.71 4.30
SC 17.10 10.01 17.10 10.23
B + SC 3.21 4.64 3.10 4.81
Table 4: Relative errors (%) for the length ratio ($L_{p},L_{n}$) estimation problem.
@@
Battery length and length ratio Finally, we extended the problem to simultaneously estimate both the battery length and the length ratio. This combined task represents a significantly more challenging problem due to the increased dimensionality and the inherent coupling between the two parameters. Accurate estimation in such scenarios requires the framework to handle dynamic changes in both domain lengths and parameter dependencies throughout the optimization process. Starting with the same initialization for $\hat{L},\hat{\rho}_{p}$, and $\hat{\rho}_{n}$, the B + SC method estimated $\hat{L}=2.971$, $\hat{L}_{p}=1.614\cdot 10^{-4}$, and $\hat{L}_{n}=0.9374\cdot 10^{-4}$. The relative errors for these estimates were $8.85\cdot 10^{-2}$, $7.23\cdot 10^{-2}$, and $6.26\cdot 10^{-2}$, respectively.
Method w/o Noise w/ Noise
$L$ $L_{p}$ $L_{n}$ $L$ $L_{p}$ $L_{n}$
Baseline 69.66 77.67 40.01 69.66 77.67 39.98
B 52.01 10.86 13.45 52.19 10.74 13.10
SC 69.51 77.57 39.16 69.51 77.57 39.13
B + SC 8.85 7.23 6.26 8.89 7.39 6.19
Table 5: Relative errors (%) for the total length and length ratio ($L,L_{p},L_{n}$) estimation problem.
The results throughout 3 different inverse problems demonstrate the robustness of the B + SC method in tackling complex coupled inverse problems.
@@ 7 Discussion and conclusion
This study has addressed the challenges of simulating Li-ion batteries using the P2D model in combination with PINNs and has proposed effective strategies to overcome these obstacles. Our methods enable reliable and accurate simulations of the P2D model, incorporating the fully nonlinear BV equation. By developing an enhanced forward solver, we successfully tackled inverse problems related to battery section lengths, achieving remarkable accuracy and computational efficiency. Despite the inherent complexity of these problems, our framework delivers results that hold significant engineering relevance for the design and optimization of Li-ion batteries. These achievements highlight the practical applicability of our approach and its ability to solve inverse problems that might not be infeasible for traditional numerical methods. However, there is still room for further development. For instance, extending our framework to incorporate multiphysics phenomena such as thermal effects or aging, which are currently excluded from our equations, presents an exciting and challenging avenue for future research. Additionally, building on our forward problem solver to develop an operator network capable of real-time predictions is another promising direction.
@@ Acknowledgements
The work of M.-S. Lee was supported by Basic Science Research Programs through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (RS-2023-00244475 and RS-2024-00462755). The work of Y. Hong was supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF), funded by the Ministry of Education (NRF-2021R1A2C1093579), the Korean government (MSIT) (RS-2023-00219980), and Hyundai Motor Company. This work was supported by the National Supercomputing Center with supercomputing resources including technical support (KSC-2023-CRE-0334).
Appendix A Additional experiments
In addition to the original 1C discharge rate which corresponds to $\tau=3500$ and $I_{\mathrm{app}}=-17.5$, we test our methods to two different discharge rates, 0.5C and 2C. Specifically, 0.5C discharge rate is $\tau=7000$ and $I_{\mathrm{app}}=-8.75$, and 2C discharge rate corresponds to $\tau=1550$ and $I_{\mathrm{app}}=-35$. Table 6 illustrates our method’s ability to simulate different discharge rates.
Discharge rate $\psi$ $c_{s}$ $\phi$ $c$
1C 4.84e-3 1.31e-2 2.31e-2 6.51e-3
0.5C 4.87e-3 2.43e-2 1.28e-2 3.58e-3
2C 1.65e-2 4.69e-2 7.15e-2 6.02e-3
Table 6: Test results (relative $L^{2}$ errors) for different discharge rates.